if matrix A =
{5 3 -2}
{4 2 1}
{7 -1 4}
a) Multiply A by the matrix
1
1
1
(i)Explain the effect on the elements of A by this product of matrices?

b) (i) Multiply 1 1 1 by A
(ii) Explain the effect on the elements of A by this product of matrices?
1
Expert's answer
2020-05-25T20:00:42-0400

Given,


"A=\begin{bmatrix}
 5 & 3&-2 \\
 4 & 2&1\\
7&-1&4
\end{bmatrix}"

a). We have to multiply to right by a column matrix

"B=\begin{bmatrix}
 1 \\
 1\\1
\end{bmatrix}"

with "A" ,thus

"\begin{bmatrix}
 5 & 3&-2 \\
 4 & 2&1\\
7&-1&4
\end{bmatrix}\begin{bmatrix}
 1 \\
 1\\1
\end{bmatrix}=\begin{bmatrix}
 5\cdot1+3\cdot1+(-2)\cdot1\\
4\cdot1+2\cdot1+1\cdot1\\
7\cdot1+ (-1)\cdot1+4\cdot1
\end{bmatrix}=\begin{bmatrix}
 6\\7\\10
\end{bmatrix}"

On multiplying "B" with"A" , the elements of "A" get multiplied by 1(multiplication of "B" is done by each rows of "A") thus there is no effect on the elements of "A" but the final result get changed completely as before multiplication "A" was "3\times3" matrix but after multiplication its a "3\times 1" column matrix.


b).Now we have to multiply transpose of "B" i.e "B^t" to left of "A" ,thus

"\begin{bmatrix}
 1 &
 1&1
\end{bmatrix}\begin{bmatrix}
 5 & 3&-2 \\
 4 & 2&1\\
7&-1&4
\end{bmatrix}=\\\begin{bmatrix}
 1\cdot 5+1\cdot4+1\cdot7&1\cdot3+1\cdot2+1\cdot(-1)&1\cdot(-2)+1\cdot1+1\cdot4
\end{bmatrix}=\\\begin{bmatrix}
 16 &
 4&3
\end{bmatrix}"

Clearly, as part (a) ,the elements of "A" get multiplied by 1(in fact this time multiplication is done by each column of "A" by "B") and the finally "A" becomes a "1\times3" row matrix.

Hence we are done.